# Daejeon 34141, South Korea

### v1

@@ 1.5. Acknowledgments
We thank Michael Kapovich for various help with geometric group theory and coarse geometry. This article began with some discussions with Michael Kapovich during the conference honoring the 60th birthday of William Goldman at the University of Maryland, College Park, in 2016. We thank Jeffrey Danciger and Florian Stecker for help with the $P$-Anosov properties, which yielded the main idea for the proof of Theorem 1.1. We also thank Herbert Abels, Richard Canary, Virginie Charette, Todd Drumm, William Goldman, François Guéritaud, Fanny Kassel, Andrés Sambarino, Nicholas Tholozan, and Konstantinos Tsouvalas for various discussions helpful for this paper. We also thank BIRS, Banff, Canada, and KIAS, Seoul, where some of this work was done. This paper is a joint work with Kapovich in the beginning and Danciger and Stecker later. Unfortunately, they did not wish to join as coauthors for various reasons.
@@ 2. Preliminary
The other purpose is to establish the Anosov properties for manifolds using coarse isometries since most of the literature works on with groups.
@@ 2.1. Convergences of geodesics
We say that a sequence of isometric rays (resp. complete isometric geodesics) $l_{i}$ converges to an isometric ray (resp. complete geodesic) $l$ (denoted as $l_{i}\to l$) if $l_{i}(t)\to l(t)$ for each $t\in[0,\infty)$ (resp. $t\in\mathbb{R}$).
@@ 2.2. Gromov hyperbolic spaces
$d_{Y}(f(x),f(x^{\prime})\leq Ld_{X}(x,x^{\prime})+C^{\prime}\hbox{ for all }x,x^{\prime}\in X.$
A map $f:X\rightarrow Y$ is an $(L,C)$-quasi-isomeric embedding if $f$ satisfies
@@
A map $f:X\rightarrow Y$ between metric spaces is called a quasi-isometry if it is a coarse Lipschitz map and admits a coarse Lipschitz coarse inverse map. (Note that these are not necessarily continuous, as we follow Druţu-Kapovich [14].)
@@
Two isometric rays $\rho_{1}$ and $\rho_{2}$ of $X$ are equivalent if $t\mapsto d(\rho_{1}(t),\rho_{2}(t))$ is bounded. This condition is equivalent to the one that their Hausdorff distance under $d$ is bounded. Assume that $X$ is $\delta$-hyperbolic. $X$ has a well-defined ideal boundary $\partial_{\infty}X$ as in Definition 3.78 of [14]; that is, the space of the equivalence classes of isometric rays. (See [9] and [20].) We denote by $\partial^{(2)}_{\infty}X=\partial_{\infty}X\times\partial_{\infty}X-\Delta$ where $\Delta$ is the diagonal set.
@@
Since $l_{1}$ and $l_{2}$ are both $(1,0)$-quasigeodesics, this follows from Proposition 3.1 of Chapter 3 of [9]. ∎
@@
Notice that $d_{X}(g(t),h(t))\leq 2|t|$ since we can follow the geodesic $g$ backward and then forward in $h$ to obtain a distance $2|t|$. Thus, the above integral is always well defined. (See Gromov [23].) Let $GX$ have this metric topology. (See Section 11.11 of [14].)
@@
There is a map $G\hat{M}\to\mathbf{UC}\hat{M}$ given by sending the complete isometric geodesic $g:\mathbb{R}\to\hat{M}$ to $(g(0),\vec{v}_{0})$ where $\vec{v}_{0}$ is the unit tangent vector at $g(0)$. This map is clearly bijective. In Section III of [34], the map $G\hat{M}\to\hat{M}$ defined by $g\mapsto g(0)$ is shown to be a quasi-isometry. Therefore, post-composing this map with a lift that sends each $g(0)$ to a vector in the fiber of $\mathbf{UC}\hat{M}$ over $g(0)$ results in a quasi-isometry. For a sequence $\{g_{i}\}$ of geodesics in $\hat{M}$, if $d_{GX}(g,g_{i})\to 0$, then $(g_{i}(0),g_{i}^{\prime}(0))\to(g(0),g^{\prime}(0))$ since otherwise, we would obtain a positive lower bound for the integral (2).
The inverse map $\mathbf{UC}\hat{M}\to G\hat{M}$ is also continuous due to the continuity of the exponential map and considerations involving (2), where $d_{X}(g(t),h(t))$ grows sublinearly. The continuity of the integral values under $g$ and $h$ can be shown by cutting off for $|t|>N$ where $N$ is independent of $g$ or $h$ by (2) and an estimate.
@@ 2.3. Flow space $G\hat{M}$ quasi-isometric to $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$
$\partial_{\infty}^{(2)}\Gamma:=\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma-\{(t,t)|t\in\partial_{\infty}\Gamma\}.$
@@
Let $\Gamma$ be a finitely generated word hyperbolic group.
@@
Champetier [8] constructs the space $G\hat{\Gamma}_{M}$ from $\hat{\Gamma}_{M}$ quasi-isometric to $\Gamma_{M}$ following Gromov. Recall from Section 2.2 that $\partial_{\infty}\hat{M}=\partial_{\infty}\Gamma_{M}$. Using the orbit map $\Gamma_{M}\to\hat{M}$ sending $\gamma$ to $\gamma(x_{0})$, we extend the map to $\hat{\Gamma}_{M}\to\hat{M}$ by sending each edge to a shortest geodesic.
Let $I:G\hat{\Gamma}_{M}\to G\hat{M}$ be a map that sends a geodesic $g$ in $\hat{\Gamma}_{M}$ to a complete isometric geodesic $g^{\prime}$ in $\hat{M}$ with the same endpoints in $\partial_{\infty}\hat{M}$, and maps $g(0)$ to the nearest point on the image of $g^{\prime}$.
The map $I^{\prime}:G\hat{M}\to G\hat{\Gamma}_{M}$ can be defined by taking a complete isometric geodesic $g$ in $\hat{M}$ to a geodesic $g^{\prime}$ in $\hat{\Gamma}_{M}$ in a similar manner, with $g(0)$ going to one of the elements of $\Gamma_{M}(x_{0})$ nearest to it.
@@ 3.1. Some background
We have a decomposition to $ad(\mathfrak{a})$-eigenspaces $\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}$. Here, $\mathfrak{g}_{0}$ is the direct sum of $z(\mathfrak{g})$ and the centralizer of $\mathfrak{a}$ in $\mathfrak{g}_{s}$. The set $\Sigma\subset\mathfrak{a}^{*}=\mathrm{Hom}_{\mathbb{R}}(\mathfrak{a},\mathbb{R})$ projects to a (possibly nonreduced) root system of $\mathfrak{a}_{s}^{*},$ and each $\alpha\in\Sigma$ is called a restricted root of $\mathfrak{a}$ in $\mathfrak{g}$. There is a subset $\Delta\subset\Sigma$ called a simple system, i.e. a subset such that any root is expressed uniquely as a linear combination of elements of $\Delta$ with coefficients all of the same sign; the elements of $\Delta$ are called simple roots. There is a subset $\Sigma^{+}\subset\Sigma$ which is the set of positive roots, i.e. roots that are nonnegative linear combinations of elements of $\Delta$, where $\Sigma=\Sigma^{+}\cup\left(-\Sigma^{+}\right)$ holds additionally.
@@
Let $G={\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$. The maximal abelian subalgebra $A_{n}$ of $\mathfrak{p}$ in this case is the subspace of diagonal matrices with entries $a_{1},\dots,a_{n}$ where $a_{1}+\dots+a_{n}=0$. Let $\log\lambda_{i}:A_{n}\rightarrow{\mathbb{R}}$ denote the projection to the $i$-th factor. $\Sigma$ consists of $\alpha_{ij}:=\log\lambda_{i}-\log\lambda_{j}$ for $i\neq j$, and $\mathfrak{sl}(n,{\mathbb{R}})_{\alpha_{ij}}=\{cE_{ij}|c\in{\mathbb{R}}\}$ where $E_{ij}$ is a matrix with entry $1$ at $(i,j)$ and zero entries everywhere else.
@@
We recall Example 2.14 of [24]: Let $G$ be $\mathrm{GL}(n,{\mathbb{R}})$ seen as a real Lie group. Its derived group is $G_{s}=D(G)=\mathrm{SL}(n,{\mathbb{R}}).$ We can take $\mathfrak{a}\subset\mathfrak{gl}(n,{\mathbb{R}})$ to be the set of real diagonal matrices of size $n\times n$. As above, for $1\leq i\leq n,$ let $\log\lambda_{i}\in\mathfrak{a}^{*}$ be the evaluation of the $i$-th diagonal entry. Then
@@
(3) $\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\ldots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\}\\ \hbox{ with }1\leq i_{1}<\cdots<i_{m}\leq n-1,$
the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the group of block upper ( resp. lower) triangular matrices in $\mathrm{GL}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\ldots,i_{m}-i_{m-1},n-i_{m}$. In particular, $P_{\Delta}$ is the group of upper triangular matrices in $\mathrm{GL}(n,({\mathbb{R}})$.
@@ 3.2. $P$-Anosov representations
Let $(P^{+},P^{-})$ be a pair of opposite parabolic subgroups of ${\mathsf{GL}}(n,{\mathbb{R}})$, and let $\mathcal{F}^{{\pm}}$ denote the flag spaces ${\mathsf{GL}}(n,{\mathbb{R}})/P_{\pm}$. Let $\chi$ denote the unique open ${\mathsf{GL}}(n,{\mathbb{R}})$-orbit of the product $\mathcal{F}^{+}\times\mathcal{F}^{-}$. The product subspace $\chi$ has two ${\mathsf{GL}}(n,{\mathbb{R}})$-invariant distributions $E^{\pm}:=T_{x_{\pm}}\mathcal{F}^{\pm}$ for $(x^{+},x^{-})\in\chi$. Let $\phi_{t}$ be given as in Theorem 2.4 as the $\mathbb{R}$}-action on $F\Gamma_{M}$ whose orbits are quasi-geodesics. We denote by $(\tau_{+},\tau_{-}):F\Gamma_{M}\rightarrow F\Gamma_{M}/{\mathbb{R}}\cong\delta_{\infty}\Gamma_{M}^{(2)}$ the maps associating to a point the end points of its $\mathbb{R}$}-orbit.
@@
A representation $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is $(P^{+},P^{-})$-Anosov if there exist continuous $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{+}$, $\xi^{-}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{-}$ such that:
@@
We call $\xi^{\pm}:\partial_{\infty}F\Gamma_{M}\rightarrow{\mathcal{F}}^{\pm}$ the Anosov maps associated with $\rho:\Gamma_{M}\rightarrow G$. (See also Section 2.5 of [24] for more details.)
@@
$\begin{split}\vec{a}(g)&:=(a_{1}(g),\dots,a_{n}(g)),\\ \log\vec{a}(g)&:=(\log a_{1}(g),\dots,\log a_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n}),\hbox{Cartan projection}\\ \log\vec{\lambda}(g)&:=(\log\lambda_{1}(g),\dots,\log\lambda_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n})\hbox{ Jordan projection}\end{split}$
@@
Let a finitely generated group $G$ have a representation $\rho:G\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$.
@@
This follows by Theorem 1.3(3) of [24] since the set $\theta$ contains $k$ and nkn-k. ∎

### v2

@@ 1.5. Acknowledgments
We thank Michael Kapovich for various help with geometric group theory and coarse geometry. This article began with some discussions with Michael Kapovich during the conference honoring the 60th birthday of William Goldman at the University of Maryland, College Park, in 2016. We thank Jeffrey Danciger and Florian Stecker for help with the $P$-Anosov properties, which yielded the main idea for the proof of Theorem 1.1. We also thank Herbert Abels, Richard Canary, Virginie Charette, Todd Drumm, William Goldman, François Guéritaud, Fanny Kassel, Andrés Sambarino, Nicholas Tholozan, and Konstantinos Tsouvalas for various discussions helpful to this paper. We also thank BIRS, Banff, Canada, and KIAS, Seoul, where some of this work was done. This paper is a joint work with Kapovich in the beginning and Danciger and Stecker later. Unfortunately, they did not wish to join as coauthors for various reasons. Also, we used OpenAI to assist us with writing and Google Gemini to find related results and papers.
@@ 2. Preliminary
The other purpose is to establish the Anosov properties for manifolds using coarse isometries since most of the literature works with groups.
@@ 2.1. Convergences of geodesics
We say that a sequence of isometric rays (resp. complete isometric geodesics) $l_{i}$ converges to an isometric ray (resp. complete geodesic) $l$ (denoted as $l_{i}\rightarrow l$) if $l_{i}(t)\rightarrow l(t)$ for each $t\in[0,\infty)$ (resp. $t\in{\mathbb{R}}$).
@@ 2.2. Gromov hyperbolic spaces
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{Y}(f(x),f(x^{\prime}))}\leq Ld_{X}(x,x^{\prime})+C^{\prime}\hbox{ for all }x,x^{\prime}\in X.$
A map $f:X\rightarrow Y$ is an $(L,C)$-quasi-isometric embedding if $f$ satisfies
@@
A map $f:X\rightarrow Y$ between metric spaces is called a quasi-isometry if it is a coarse Lipschitz map that admits a coarse Lipschitz coarse inverse map. (Note that these are not necessarily continuous, as we follow Druţu-Kapovich [14].)
@@
Two isometric rays $\rho_{1}$ and $\rho_{2}$ of $X$ are equivalent if $t\mapsto d(\rho_{1}(t),\rho_{2}(t))$ is bounded. This condition is equivalent to the one that their Hausdorff distance under $d$ is bounded. Assume that $X$ is $\delta$-hyperbolic. $X$ has a well-defined ideal boundary $\partial_{\infty}X$ as in Definition 3.78 of [14]; that is, the space of the equivalence classes of isometric rays. (See [9] and [20].) We denote by $\partial^{(2)}_{\infty}X=\partial_{\infty}X\times\partial_{\infty}X\setminus\Delta$ where $\Delta$ is the diagonal set.
@@
Since $l_{1}$ and $l_{2}$ are both $(1,0)$-quasi-geodesics, this follows from Proposition 3.1 of Chapter 3 of [9]. ∎
@@
Notice that $d_{X}(g(t),h(t))\leq 2|t|+C$ for a constant $C$ since we can follow the geodesic $g$ backward and then forward in $h$ to obtain a distance $2|t|$. Thus, the above integral is always well defined. (See Gromov [23].) Let $GX$ have this metric topology. (See Section 11.11 of [14].)
@@
There is a map $G\hat{M}\rightarrow\mathbf{UC}\hat{M}$ given by sending each complete isometric geodesic $g:{\mathbb{R}}\rightarrow\hat{M}$ to $(g(0),\vec{v}_{0})$ where $\vec{v}_{0}$ is the unit tangent vector at $g(0)$. This map is clearly bijective. In Section III of [35], the map $G\hat{M}\rightarrow\hat{M}$ defined by $g\mapsto g(0)$ is shown to be a quasi-isometry. Therefore, post-composing this map with a lift that sends each $g(0)$ to a vector in the fiber of $\mathbf{UC}\hat{M}$ over $g(0)$ results in a quasi-isometry. For a sequence $\{g_{i}\}$ of geodesics in $\hat{M}$, if $d_{GX}(g,g_{i})\to 0$, then $(g_{i}(0),g_{i}^{\prime}(0))\to(g(0),g^{\prime}(0))$ since otherwise, we would obtain a positive lower bound for the integral (2).
The inverse map $\mathbf{UC}\hat{M}\rightarrow G\hat{M}$ is also continuous due to the continuity of the exponential map and considerations involving (2), where $d_{X}(g(t),h(t))$ grows sublinearly. The continuity of the integral values under $g$ and $h$ can be shown by cutting off for $|t|>N$ where $N$ is independent of $g$ or $h$ by (2) and an estimate.
@@ 2.3. Flow space $G\hat{M}$ quasi-isometric to $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$
$\partial_{\infty}^{(2)}\Gamma:=\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma\setminus\{(t,t)\mid t\in\partial_{\infty}\Gamma\}.$
@@
Let $\Gamma$ be a finitely generated word-hyperbolic group.
@@
Champetier [8] constructs the space $G\hat{\Gamma}_{M}$ from $\hat{\Gamma}_{M}$ quasi-isometric to $\Gamma_{M}$ following Gromov. Recall from Section 2.2 that $\partial_{\infty}\hat{M}=\partial_{\infty}\Gamma_{M}$. Using the orbit map $\Gamma_{M}\rightarrow\hat{M}$ sending $\gamma$ to $\gamma(x_{0})$, we extend the map to $\hat{\Gamma}_{M}\rightarrow\hat{M}$ by sending each edge to a shortest geodesic.
Let $I:G\hat{\Gamma}_{M}\rightarrow G\hat{M}$ be a map that sends a geodesic $g$ in $\hat{\Gamma}_{M}$ to a complete isometric geodesic $g^{\prime}$ in $\hat{M}$ with the same endpoints in $\partial_{\infty}\hat{M}$, and maps $g(0)$ to the nearest point on the image of $g^{\prime}$.
The map $I^{\prime}:G\hat{M}\rightarrow G\hat{\Gamma}_{M}$ can be defined by taking a complete isometric geodesic $g$ in $\hat{M}$ to a geodesic $g^{\prime}$ in $\hat{\Gamma}_{M}$ in a similar manner, with $g(0)$ going to one of the elements of $\Gamma_{M}(x_{0})$ nearest to it.
@@ 3.1. Some background
We have a decomposition to $ad(\mathfrak{a})$-eigenspaces $\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}$. Here, $\mathfrak{g}_{0}$ is the direct sum of $z(\mathfrak{g})$ and the centralizer of $\mathfrak{a}$ in $\mathfrak{g}_{s}$. The set $\Sigma\subset\mathfrak{a}^{*}={\rm Hom}_{{\mathbb{R}}}(\mathfrak{a},{\mathbb{R}})$ projects to a (possibly nonreduced) root system of $\mathfrak{a}_{s}^{*},$ and each $\alpha\in\Sigma$ is called a restricted root of $\mathfrak{a}$ in $\mathfrak{g}$. There is a subset $\Delta\subset\Sigma$ called a simple system, i.e. a subset such that any root is expressed uniquely as a linear combination of elements of $\Delta$ with coefficients all of the same sign; the elements of $\Delta$ are called simple roots. There is a subset $\Sigma^{+}\subset\Sigma$ which is the set of positive roots, i.e. roots that are nonnegative linear combinations of elements of $\Delta$, where $\Sigma=\Sigma^{+}\cup\left(-\Sigma^{+}\right)$ holds additionally.
@@
Let $G={\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$. The maximal abelian subalgebra $A_{n}$ of $\mathfrak{p}$ in this case is the subspace of diagonal matrices with entries $a_{1},\dots,a_{n}$ where $a_{1}+\dots+a_{n}=0$. Let $\log\lambda_{i}:A_{n}\rightarrow{\mathbb{R}}$ denote the projection to the $i$-th factor. $\Sigma$ consists of $\alpha_{ij}:=\log\lambda_{i}-\log\lambda_{j}$ for $i\neq j$, and the root space $\mathfrak{sl}(n,{\mathbb{R}})_{\alpha_{ij}}=\{cE_{ij}|c\in{\mathbb{R}}\}$ where $E_{ij}$ is a matrix with an entry $1$ at $(i,j)$ and zero entries everywhere else.
@@
We recall Example 2.14 of [24]: Let $G$ be ${\mathsf{GL}}(n,{\mathbb{R}})$ seen as a real Lie group. Its derived group is $G_{s}=D(G)={\mathsf{SL}}(n,{\mathbb{R}}).$ We can take $\mathfrak{a}\subset\mathfrak{gl}(n,{\mathbb{R}})$ to be the set of real diagonal matrices of size $n\times n$. As above, for $1\leq i\leq n,$ let $\log\lambda_{i}\in\mathfrak{a}^{*}$ be the evaluation of the $i$-th diagonal entry. Then
@@
(3) $\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\dots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\}\\ \hbox{ with }1\leq i_{1}<\cdots<i_{m}\leq n-1,$
the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the group of block upper ( resp. lower) triangular matrices in ${\mathsf{GL}}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\dots,i_{m}-i_{m-1},n-i_{m}$. In particular, $P_{\Delta}$ is the group of upper triangular matrices in ${\mathsf{GL}}(n,{\mathbb{R}})$.
@@ 3.2. $P$-Anosov representations
Let $(P^{+},P^{-})$ be a pair of opposite parabolic subgroups of ${\mathsf{GL}}(n,{\mathbb{R}})$, and let $\mathcal{F}^{{\pm}}$ denote the flag spaces ${\mathsf{GL}}(n,{\mathbb{R}})/P_{\pm}$. Let $\chi$ denote the unique open ${\mathsf{GL}}(n,{\mathbb{R}})$-orbit of the product $\mathcal{F}^{+}\times\mathcal{F}^{-}$. The product subspace $\chi$ has two ${\mathsf{GL}}(n,{\mathbb{R}})$-invariant distributions $E^{\pm}:=T_{x_{\pm}}\mathcal{F}^{\pm}$ for $(x^{+},x^{-})\in\chi$. Suppose that $\Gamma_{M}$ is a finitely generated word-hyperbolic group. Let $\phi_{t}$ be given as in Theorem 2.4 as the $\mathbb{R}$}-action on $F\Gamma_{M}$ whose orbits are quasi-geodesics. We denote by $(\tau_{+},\tau_{-}):F\Gamma_{M}\rightarrow F\Gamma_{M}/{\mathbb{R}}\cong\partial_{\infty}\Gamma_{M}^{(2)}$ the maps associating to a point the endpoints of its $\mathbb{R}$}-orbit.
@@
Suppose that $\Gamma_{M}$ is a finitely generated word-hyperbolic group. A representation $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is $(P^{+},P^{-})$-Anosov if there exist continuous $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{+}$, $\xi^{-}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{-}$ such that:
@@
We call $\xi^{\pm}:\partial_{\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{M}}\rightarrow$$\mathcal{F}$$^{\pm}$ the Anosov maps associated with $\rho:\Gamma_{M}\rightarrow G$. (See also Section 2.5 of [24] for more details.)
@@
$\begin{split}\vec{a}(g)&:=(a_{1}(g),\dots,a_{n}(g)),\\ \log\vec{a}(g)&:=(\log a_{1}(g),\dots,\log a_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hbox{(Cartan projection)}},\\ \log\vec{\lambda}(g)&:=(\log\lambda_{1}(g),\dots,\log\lambda_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hbox{(Jordan projection)}}\end{split}$
@@
Let a finitely generated word-hyperbolic group $G$ have a representation $\rho:G\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$.
@@
This follows from Theorem 1.3(3) of [24] since the set $\theta$ contains $k$ and nkn-k. Or one can use Propositions 4.5 and 4.9 of [5].